There were 24 students in Mrs.Terry's class. 1/6 of the students were born in the first 4 months of the year. 1/2 of the remaining students were born in the months of May through Augest.There are no September birthdays, but 1/5 of the remaining birthdays are in October. If 1 student was born in November, how many December birthdays are there? Show your work.
Answer by rwm(914)

total class =24
1/6 of 24 =4
24-4=20 remaining for may to december
1/2 *20=10
24-4-10=10 remaining for september to december
1/5*10=2
we now have 8 remaining for november and december
8-1=7 remaining for december
7 were born in December